Question 9 0 out of 5 points Problem 3C: The best alternative using B/C ratio analysis is (5 points - Justify your answer wit data for four nutually exlcusive alternatives are given in the table below, Assume a life of 7 years and a MARR of 9%
Answer #1

Correct Answer:

Alt. C

Working note:

For, alternative A, equivalent annual cost = 5600*(A/P, 9%, 7) = 5600*.1987

For, alternative A, equivalent annual cost = $1112.72

So,

B/C ratio of alternative A = 1400/1112.72 = 1.26

For, alternative B, equivalent annual cost = 3400*(A/P, 9%, 7) = 3400*.1987

For, alternative B, equivalent annual cost = $675.58

B/C ratio of alternative B = 1000/675.58 = 1.48

For, alternative C, equivalent annual cost = 1200*(A/P, 9%, 7) = 1200*.1987

For, alternative C, equivalent annual cost = $238.44

B/C ratio of alternative C = 400/238.44 = 1.68

Since alternative C has the highest B/C ratio, then alternative C should be selected.